OpenStudy (loser66):

y^(6) +y =0 I mean the 6th derivative Please help

OpenStudy (mendicant_bias):

What do you specifically need to find?

OpenStudy (loser66):

the general solution for ODE

OpenStudy (mendicant_bias):

I don't know what ODE is, never heard the term before.

OpenStudy (loser66):

characteristics equation for that is \[\lambda ^6 +1=0\]

OpenStudy (mendicant_bias):

Nevermind, this is over my head. I haven't taken Calc II, III, or PDE. Good luck.

OpenStudy (loser66):

\[lambda^6 = -1\rightarrow \lambda= \pm i \]3 times, but cannot get the answer.

OpenStudy (anonymous):

u need to find 6th roots of \(-1\) but there is a neater way to do this\[\lambda^6+1=0\]\[\lambda^6-i^6=0\]\[(\lambda^3-i^3)(\lambda^3+i^3)=0\]\[(\lambda-i)(\lambda^2+\lambda i-1)(\lambda+i)(\lambda^2-\lambda i-1)=0\]

OpenStudy (loser66):

oh yeah, you are right! but what if I steer in this way \[-1= e^{i(\pi+2m\pi)}\]

OpenStudy (anonymous):

thats very good also...

OpenStudy (loser66):

hey, not good at all. I did and got stuck. hehe. but from yours, how to solve the second and the last one to get lambda?

OpenStudy (anonymous):

those are quadratics, use the formula :)

OpenStudy (loser66):

thank you, let me try. however, I still have a question: I know you make 1 = - (-1) = - (i^2) but don't get how to make it = i^6

OpenStudy (anonymous):

well \(i^4=1\) u can put \(i^4\) there

OpenStudy (loser66):

ok, i^4*i^2 =i^6= -1? or i^2 = -1 and (i^2)^3= (-1)^3 = -1? right?

OpenStudy (anonymous):

right

OpenStudy (loser66):

bingo, since your smartscore is 99, no need to give you medal , right? just "thanks a ton" XD

OpenStudy (anonymous):

:) thats right :D very welcome

OpenStudy (anonymous):

hey u made me champion with that medal :D :) :) thanks

OpenStudy (loser66):

hahahaa.... so that means you are champion now? congratulation

OpenStudy (anonymous):

lol, yes, thanks

OpenStudy (loser66):

yw